# STT-GS: Sample-Then-Transmit Edge Gaussian Splatting with Joint Client Selection and Power Control

### v3

@@ I Introduction
Reconstructing three-dimensional (3D) environments is crucial for embodied robotics and low-altitude economy (LAE) applications [agarwal2011building, zhu2018very, fruh2004automated, xu20193d, li2024seamless]. However, conventional 3D reconstruction methods rely on neural radiance fields (NeRF), which often incur considerable computational overhead [mildenhall2021nerf]. To address this limitation, 3D Gaussian splatting (GS)[kerbl20233d] has recently been proposed through providing enhanced computational efficiency through explicit 3D geometric representations. For instance, 3D GS can be trained within ten minutes and is capable of rendering images at over 50 Hz [jena2025sparfels]. In practice, reconstructing large-scale scenes inevitably requires leveraging data distributed across multiple robots [maboudi2023review]. This motivates the adoption of the edge GS (EGS) paradigm, which aggregates data from distributed clients and trains a global GS model at the edge server.
@@ I-A Related Work
Generally, edge resource optimization relies heavily on the designed objective functions. Conventional approaches [yu2004iterative, zhang2024efficient, schubert2004solution, al2011achieving, beck20141, wang2020learning, wang2020machine, yoo2019learning] adopt communication throughput [yu2004iterative, zhang2024efficient, schubert2004solution], user fairness [zheng2016wireless], power consumption [schubert2004solution, zhang2024integrated], or general-purpose learning performance [wang2020learning, wang2020machine, yoo2019learning, luo2023joint, luo2024deep] as their optimization objectives. Nonetheless, these metrics are not tailored for GS and cannot explicitly maximize the actual rendering quality. Recent studies attempt to design loss functions tailored to visual tasks, such as adjusting transmission priorities via feature matching errors [wang2024mba]. However, these methods fail to account for the non-uniform contribution of Gaussian primitives to the eventual rendering quality. There also exist distributed GS methods that account for the rendering loss based on gradient importance [suzuki2024fed3dgs]. Yet, such importance evaluation is often based on heuristics (e.g., gradient norm thresholding) and cannot fully reflect the geometric data characteristics. Therefore, how to define a proper objective function for EGS still remains an open question.
Given a certain objective function, the next step is to conduct optimization. Conventional edge resource optimization mainly focuses on physical-layer aspects. For instance, the learning centric power allocation method is proposed in [wang2020learning], and the multi-antenna beamforming design is proposed in [liang2024communication] for edge intelligence systems. To achieve the best rendering performance, there is a paradigm shift towards cross-layer optimization [marchisio2019deep, zhou2019edge, feng2024edge, duan2023combining]. In this direction, prior works focus on edge video streaming [feng2024edge] or edge machine learning [zhou2019edge, duan2023combining], improving general-purpose multimedia or learning performance by jointly optimizing physical-layer (e.g., power allocation) and application-layer (e.g., task scheduling) parameters. Nevertheless, these approaches neglect the GS rendering requirements. Indeed, the clients should be prioritized according to their potential contributions to multi-view GS rendering, while taking their channel conditions into account. Such cross-layer GS optimization has not been addressed in existing literature.
@@ I-B Contributions of This Work
To fill the research gap, this paper formulates a novel GS-oriented objective function that distinguishes the heterogeneous view contributions of different clients. This objective function, which is built upon the uncertainty sampling theory [yoo2019learning, zhu2009active, beluch2018power] and the vanilla GS loss function [kerbl20233d], enables the edge server to maximize the information gain brought by the collected GS data. Subsequently, we integrate the new objective function with the communication conditions for effective GS cross-layer optimization. This leads to a joint client selection and power control (JCSPC) framework, which unifies viewpoint contribution, interference mitigation, and resource allocation into a single optimization formulation.
@@
However, the above GS problem involves two key technical challenges. First, the GS-oriented objective, being a function of clients’ images, is inaccessible prior to data transmissions, while data transmissions conversely require the guidance from the GS-oriented objective. This thereby introduces a so-called causality dilemma (i.e., chicken-and-egg paradox). Second, JCSPC involves nonlinear coupling among binary client selection variables and continuous power control variables, due to potential cross-layer multi-user interference. Consequently, it belongs to a nonconvex mixed integer nonlinear programming (NMINLP) problem, which is nonconvex even after continuous relaxation. Existing approaches, e.g., convex optimization [diamond2016cvxpy, yu2004iterative, schubert2004solution] or majorization minimization ($M$) [sun2016majorization, wang2020learning, wen2023byzantine, wen2024augment], are not applicable.
@@
To address the first challenge, this paper proposes a sample-then-transmit EGS (STT-GS for short) strategy, which first samples a subset of images as pilot data from each client for loss prediction and then prioritizes resources towards more valuable clients based on the first-stage evaluation. Since sampling also involves communication costs, it is necessary to improve the sampling efficiency and reduce the pilot overhead. Thus, we propose feature-domain clustering (FDC) to select the most representative data, where the sampling ratio of FDC is determined using cross-validation. Subsequently, we propose a fast iterative bisection search algorithm for pilot transmission time minimization (PTTM). To tackle the second challenge, we propose a penalty alternating majorization minimization (PAMM) method to solve the JCSPC problem. Our PAMM first adopts variable splitting to decouple the selection and power variables involved in constraints, and leverages penalization to ensure constraint feasibility. Such reformulation yields a separable structure that can be safely handled by alternating minimization (AM), which is guaranteed to converge to a local minimum. [beck20141]. By further incorporating $M$ into AM for joint optimization, the proposed PAMM algorithm effectively solves the NMINLP problem with polynomial computational complexity.
@@
We evaluate the proposed STT-GS scheme along with the FDC, PTTM, PAMM algorithms exploiting two real-world drone gathered LAE datasets, i.e., rubble-pixsfm and building-pixsfm [turki2022mega]. Experimental results show that the proposed scheme significantly enhances the rendering quality and strictly satisfies the communication resource budget. In particular, the proposed STT-GS outperforms the MaxRate [zhang2024efficient] and Fairness [zheng2016wireless] schemes by 4.50% and 7.81% in terms of peak signal-to-noise ratio (PSNR). Furthermore, through ablation studies, we also confirm the indispensability of FDC, PTTM, and PAMM algorithms. To the best of our knowledge, this is the first attempt to integrate GS features and communication constraints into a unified framework.
@@ II Edge Gaussian Splatting
We consider an EGS system as shown in Fig. 1, which consists of an edge server with NN antennas and $K$ single-antenna clients. The objective of the EGS system is to reconstruct a 3D scene by aggregating distributed data from clients $\mathcal{K}=\{1,\cdots,K\}$ and training a global GS model $\mathcal{S}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots\}$ using edge GPUs, where each $\mathcal{G}_{i}$ denotes a 3D Gaussian with trainable parameters. Below, we present our system model and problem formulation in detail.
@@ II-A System Model
where $\mathbf{v}_{i,k}\in\mathbb{R}^{3LW}$ denotes the $i$-th image of client $k$, where $L$ and $W$ are the length and width of camera images respectively, and the coefficient 33 accounts for the red green blue (RGB) channels. The $\mathbf{s}_{i,k}\in\mathbb{R}^{6}$ represents the associated 6D camera pose [kerbl20233d], and $l_{k}=|\mathcal{D}_{k}|$ is the number of samples at client $k$. The data volume in bits of each sample is $V_{k}$. These data can be locally generated at each client by adopting monocular cameras and localization packages along the trajectory $(\mathbf{s}_{1,k},\mathbf{s}_{2,k},\cdots,\mathbf{s}_{l_{k},k})$, which are standard modules equipped at modern robot platforms [xu2022fast] (e.g., drones).
To avoid multi-user interference during dataset uploading, a client selection module with a binary vector $\mathbf{x}=[x_{1},\cdots,x_{K}]^{T}$ and $x_{k}\in\{0,1\}$ is introduced, where $x_{k}=1$ indicates that the $k$-th client is being selected and $x_{k}=0$ otherwise. For a selected client $k$, e.g., $x_{k}=1$, it transmits its signal $z_{k}$ with power $\mathbb{E}[|z_{k}|^{2}]=p_{k}$ for uploading $\mathcal{D}_{k}$. Accordingly, the aggregated signal [wang2020learning] received at the edge server is written as
@@
where $\mathbf{r}=[r_{1},\cdots,r_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is the received signal, $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ denotes the channel from the $k$-th client to the edge server, and $\mathbf{n}\in\mathbb{C}^{N\times 1}$ is additive white Gaussian noise (AWGN) with zero mean and covariance $\mathbb{E}\{\mathbf{n}\mathbf{n}^{H}\}=\sigma^{2}\mathbf{I}_{N}$, with $\mathbf{I}_{N}$ being the identity matrix of size $N\times N$. The wireless channel is assumed to follow Rician fading[wang2020angle], i.e.,
@@
To separate the useful components $\mathbf{z}=[z_{1},...,z_{k}]$ from the received signal $\mathbf{r}$, it is necessary to mitigate the interference and noise. An ideal approach is to apply a minimum mean square error (MMSE) combing vector to process the signal $\mathbf{r}$ [lu2014overview]. However, MMSE involves calculating the matrix inverse, which is a computationally intensive operation especially for EGS with a large number of antennas. Therefore, we choose the celebrated maximum ratio combining (MRC) as in [song2025distributed, wang2020learning] with $\mathbf{w}_{k}=\left\|\mathbf{h}_{k}\right\|_{2}^{-1}\mathbf{h}_{k}$ due to its high computational efficiency. It has been proved in [song2025distributed] that MRC achieves an asymptotically equivalent performance compared to MMSE (i.e., the benefit brought by MMSE over MRC vanishes as NN increases).
@@ II-B Problem Formulation
In the considered system, the design variables that can be controlled are the client selection vector $\mathbf{x}=[x_{1},\cdots,x_{K}]^{T}$ and transmit power vector $\mathbf{p}=[p_{1},\cdots,p_{K}]^{T}$. To ensure low power consumption and limited inter-user interference, both the individual and total powers must satisfy their prescribed limits, given by $P_{\text{max}}$ and $P_{\mathrm{sum}}$, respectively. Specifically, each client’s transmit power must fulfill the constraint $\{0\leq p_{k}\leq P_{\text{max}},\forall k\}$ and $\sum_{k=1}^{K}x_{k}p_{k}\leq P_{\mathrm{sum}}$ [yu2004iterative, zhang2024efficient, al2011achieving, beck20141]. On the other hand, the dataset uploading at each selected client must be completed within a time threshold $T$, i.e., $x_{k}V_{k}|\mathcal{D}_{k}|R_{k}^{-1}\leq T$. [^1] One may wonder if the data uploading time $T$ is negligible compared to the GS model training time. This is in fact, not true, due to the recent advancements in 3D GS training algorithms [mallick2024taming, chen2025dashgaussian] (e.g., DashGaussian [chen2025dashgaussian]), which have dramatically reduced model training time to merely $3.23$ minutes. Given this significant time reduction in model training, introducing a time threshold $T$ for data transmission becomes essential for EGS.
Having the power and time constraints satisfied, it is then crucial to maximize the information gain contributed by $\mathcal{E}\setminus\mathcal{E}^{\prime}=\bigcup_{k\in\mathcal{X}}\mathcal{D}_{k}$ for better 3D reconstruction. In other words, we need to collect those datasets that could change the GS model from $\mathcal{S}^{{}^{\prime}}$ to $\mathcal{S}$ with the maximum extent. According to the uncertainty sampling theory [yoo2019learning, zhu2009active, beluch2018power], this is equivalent to maximizing the total inference loss of model $\mathcal{S}^{{}^{\prime}}$ on $\bigcup_{k\in\mathcal{X}}\mathcal{D}_{k}$, denoted as $C(\mathcal{S}^{{}^{\prime}},\{\mathcal{D}_{k}\},\mathbf{x})$. Specifically, the GS inference function $\mathcal{R}(\cdot|\mathcal{S}^{{}^{\prime}})$ is able to render a photo-realistic image from a camera pose as $\widehat{\mathbf{v}}_{i,k}=\mathcal{R}(\mathbf{s}_{i,k}|\mathcal{S}^{{}^{\prime}})$. To derive an explicit form of $C$, we need to measure the difference between the rendered image $\widehat{\mathbf{v}}_{i,k}$ (produced using $\mathcal{S}^{{}^{\prime}}$) and ground-truth image $\mathbf{v}_{i,k}$. According to [kerbl20233d], a proper metric is the vanilla GS that adopts the following loss function:
@@
respectively, with $\mathcal{L}_{\mathsf{SSIM}}$ being the SSIM function detailed in [wang2011ssim, Eqn. 5], and the weight $\lambda\in[0,1]$ is set to $\lambda=0.2$ according to [kerbl20233d].
@@ III Proposed Sample-Then-Transmit Strategy
The above 88 steps can be categorized into two main stages, i.e., EGS loss prediction and EGS full transmission, where each stage consists of 44 steps.
@@ IV EGS Loss Prediction
Given $\rho_{k}$, the next question is how to determine $\tilde{\mathcal{D}}_{k}$. A naive approach is to adopt random sampling, which randomly chooses an image inside $\mathcal{D}_{k}$ each time until $\lvert\tilde{\mathcal{D}}_{k}\rvert$ images are collected [mayer2020adversarial]. Another approach is to adopt uniform sampling, which orders the images in $\mathcal{D}_{k}$ according to the timestamps and selects an image for transmission for every $\lvert\mathcal{D}_{k}\rvert/\lvert\tilde{\mathcal{D}}_{k}\rvert$ image frames. These methods do not take the importance of images into account, and may lead to redundancy in transmitting similar images. For example, if the robot remains stationary, multiple images may be nearly identical, uploading them yields little additional value. To address the problem, the following subsection will present a method that improves sampling efficiency compared to the above approaches.

### v4

@@ I Introduction
Reconstructing three-dimensional (3D) environments is crucial for embodied robotics and low-altitude economy (LAE) applications [1, 2, 3, 4, 5]. However, conventional 3D reconstruction methods rely on neural radiance fields (NeRF), which often incur considerable computational overhead [6]. To address this limitation, 3D Gaussian splatting (GS)[7] has recently been proposed through providing enhanced computational efficiency through explicit 3D geometric representations. For instance, 3D GS can be trained within ten minutes and is capable of rendering images at over 50 Hz [8]. In practice, reconstructing large-scale scenes inevitably requires leveraging data distributed across multiple robots [9]. This motivates the adoption of the edge GS (EGS) paradigm, which aggregates data from distributed clients and trains a global GS model at the edge server.
@@ I-A Related Work
Generally, edge resource optimization relies heavily on the designed objective functions. Conventional approaches [10, 11, 12, 13, 14, 15, 16, 17] adopt communication throughput [10, 11, 12], user fairness [18], power consumption [12, 19], or general-purpose learning performance [15, 16, 17, 20, 21] as their optimization objectives. Nonetheless, these metrics are not tailored for GS and cannot explicitly maximize the actual rendering quality. Recent studies attempt to design loss functions tailored to visual tasks, such as adjusting transmission priorities via feature matching errors [22]. However, these methods fail to account for the non-uniform contribution of Gaussian primitives to the eventual rendering quality. There also exist distributed GS methods that account for the rendering loss based on gradient importance [23]. Yet, such importance evaluation is often based on heuristics (e.g., gradient norm thresholding) and cannot fully reflect the geometric data characteristics. Therefore, how to define a proper objective function for EGS still remains an open question.
Given a certain objective function, the next step is to conduct optimization. Conventional edge resource optimization mainly focuses on physical-layer aspects. For instance, the learning centric power allocation method is proposed in [15], and the multi-antenna beamforming design is proposed in [24] for edge intelligence systems. To achieve the best rendering performance, there is a paradigm shift towards cross-layer optimization [25, 26, 27, 28]. In this direction, prior works focus on edge video streaming [27] or edge machine learning [26, 28], improving general-purpose multimedia or learning performance by jointly optimizing physical-layer (e.g., power allocation) and application-layer (e.g., task scheduling) parameters. Nevertheless, these approaches neglect the GS rendering requirements. Indeed, the clients should be prioritized according to their potential contributions to multi-view GS rendering, while taking their channel conditions into account. Such cross-layer GS optimization has not been addressed in existing literature.
@@ I-B Contributions of This Work
To fill the research gap, this paper formulates a novel GS-oriented objective function that distinguishes the heterogeneous view contributions of different clients. This objective function, which is built upon the uncertainty sampling theory [17, 29, 30] and the vanilla GS loss function [7], enables the edge server to maximize the information gain brought by the collected GS data. Subsequently, we integrate the new objective function with the communication conditions for effective GS cross-layer optimization. This leads to a joint client selection and power control (JCSPC) framework, which unifies viewpoint contribution, interference mitigation, and resource allocation into a single optimization formulation.
@@
However, the above GS problem involves two key technical challenges. First, the GS-oriented objective, being a function of clients’ images, is inaccessible prior to data transmissions, while data transmissions conversely require the guidance from the GS-oriented objective. This thereby introduces a so-called causality dilemma (i.e., chicken-and-egg paradox). Second, JCSPC involves nonlinear coupling among binary client selection variables and continuous power control variables, due to potential cross-layer multi-user interference. Consequently, it belongs to a nonconvex mixed integer nonlinear programming (NMINLP) problem, which is nonconvex even after continuous relaxation. Existing approaches, e.g., convex optimization [31, 10, 12] or majorization minimization (MM) [32, 15, 33, 34], are not applicable.
@@
To address the first challenge, this paper proposes a sample-then-transmit EGS (STT-GS for short) strategy, which first samples a subset of images as pilot data from each client for loss prediction and then prioritizes resources towards more valuable clients based on the first-stage evaluation. Since sampling also involves communication costs, it is necessary to improve the sampling efficiency and reduce the pilot overhead. Thus, we propose feature-domain clustering (FDC) to select the most representative data, where the sampling ratio of FDC is determined using cross-validation. Subsequently, we propose a fast iterative bisection search algorithm for pilot transmission time minimization (PTTM). To tackle the second challenge, we propose a penalty alternating majorization minimization (PAMM) method to solve the JCSPC problem. Our PAMM first adopts variable splitting to decouple the selection and power variables involved in constraints, and leverages penalization to ensure constraint feasibility. Such reformulation yields a separable structure that can be safely handled by alternating minimization (AM), which is guaranteed to converge to a local minimum. [14]. By further incorporating MM into AM for joint optimization, the proposed PAMM algorithm effectively solves the NMINLP problem with polynomial computational complexity.
@@
We evaluate the proposed STT-GS scheme along with the FDC, PTTM, PAMM algorithms exploiting two real-world drone gathered LAE datasets, i.e., rubble-pixsfm and building-pixsfm [35]. Experimental results show that the proposed scheme significantly enhances the rendering quality and strictly satisfies the communication resource budget. In particular, the proposed STT-GS outperforms the MaxRate [11] and Fairness [18] schemes by 4.50% and 7.81% in terms of peak signal-to-noise ratio (PSNR). Furthermore, through ablation studies, we also confirm the indispensability of FDC, PTTM, and PAMM algorithms. To the best of our knowledge, this is the first attempt to integrate GS features and communication constraints into a unified framework.
@@ II Edge Gaussian Splatting
We consider an EGS system as shown in Fig. 1, which consists of an edge server with $N$ antennas and $K$ single-antenna clients. The objective of the EGS system is to reconstruct a 3D scene by aggregating distributed data from clients $\mathcal{K}=\{1,\cdots,K\}$ and training a global GS model $\mathcal{S}=\{\mathcal{G}_{1},\mathcal{G}_{2},\cdots\}$ using edge GPUs, where each $\mathcal{G}_{i}$ denotes a 3D Gaussian with trainable parameters. Below, we present our system model and problem formulation in detail.
@@ II-A System Model
where $\mathbf{v}_{i,k}\in\mathbb{R}^{3LW}$ denotes the $i$-th image of client $k$, where $L$ and $W$ are the length and width of camera images respectively, and the coefficient $3$ accounts for the red green blue (RGB) channels. The $\mathbf{s}_{i,k}\in\mathbb{R}^{6}$ represents the associated 6D camera pose [7], and $l_{k}=|\mathcal{D}_{k}|$ is the number of samples at client $k$. The data volume in bits of each sample is $V_{k}$. These data can be locally generated at each client by adopting monocular cameras and localization packages along the trajectory $(\mathbf{s}_{1,k},\mathbf{s}_{2,k},\cdots,\mathbf{s}_{l_{k},k})$, which are standard modules equipped at modern robot platforms [36] (e.g., drones).
To avoid multi-user interference during dataset uploading, a client selection module with a binary vector $\mathbf{x}=[x_{1},\cdots,x_{K}]^{T}$ and $x_{k}\in\{0,1\}$ is introduced, where $x_{k}=1$ indicates that the $k$-th client is being selected and $x_{k}=0$ otherwise. For a selected client $k$, e.g., $x_{k}=1$, it transmits its signal $z_{k}$ with power $\mathbb{E}[|z_{k}|^{2}]=p_{k}$ for uploading $\mathcal{D}_{k}$. Accordingly, the aggregated signal [15] received at the edge server is written as
@@
where $\mathbf{r}=[r_{1},\cdots,r_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is the received signal, $\mathbf{h}_{k}\in\mathbb{C}^{N\times 1}$ denotes the channel from the $k$-th client to the edge server, and $\mathbf{n}\in\mathbb{C}^{N\times 1}$ is additive white Gaussian noise (AWGN) with zero mean and covariance $\mathbb{E}\{\mathbf{n}\mathbf{n}^{H}\}=\sigma^{2}\mathbf{I}_{N}$, with $\mathbf{I}_{N}$ being the identity matrix of size $N\times N$. The wireless channel is assumed to follow Rician fading[37], i.e.,
@@
To separate the useful components $\mathbf{z}=[z_{1},...,z_{k}]$ from the received signal $\mathbf{r}$, it is necessary to mitigate the interference and noise. An ideal approach is to apply a minimum mean square error (MMSE) combing vector to process the signal $\mathbf{r}$ [38]. However, MMSE involves calculating the matrix inverse, which is a computationally intensive operation especially for EGS with a large number of antennas. Therefore, we choose the celebrated maximum ratio combining (MRC) as in [39, 15] with $\mathbf{w}_{k}=\left\|\mathbf{h}_{k}\right\|_{2}^{-1}\mathbf{h}_{k}$ due to its high computational efficiency. It has been proved in [39] that MRC achieves an asymptotically equivalent performance compared to MMSE (i.e., the benefit brought by MMSE over MRC vanishes as $N$ increases).
@@ II-B Problem Formulation
In the considered system, the design variables that can be controlled are the client selection vector $\mathbf{x}=[x_{1},\cdots,x_{K}]^{T}$ and transmit power vector $\mathbf{p}=[p_{1},\cdots,p_{K}]^{T}$. To ensure low power consumption and limited inter-user interference, both the individual and total powers must satisfy their prescribed limits, given by $P_{\text{max}}$ and $P_{\mathrm{sum}}$, respectively. Specifically, each client’s transmit power must fulfill the constraint $\{0\leq p_{k}\leq P_{\text{max}},\forall k\}$ and $\sum_{k=1}^{K}x_{k}p_{k}\leq P_{\mathrm{sum}}$ [10, 11, 13, 14]. On the other hand, the dataset uploading at each selected client must be completed within a time threshold $T$, i.e., $x_{k}V_{k}|\mathcal{D}_{k}|R_{k}^{-1}\leq T$. [^1] One may wonder if the data uploading time $T$ is negligible compared to the GS model training time. This is in fact, not true, due to the recent advancements in 3D GS training algorithms [40, 41] (e.g., DashGaussian [41]), which have dramatically reduced model training time to merely $3.23$ minutes. Given this significant time reduction in model training, introducing a time threshold $T$ for data transmission becomes essential for EGS.
Having the power and time constraints satisfied, it is then crucial to maximize the information gain contributed by $\mathcal{E}\setminus\mathcal{E}^{\prime}=\bigcup_{k\in\mathcal{X}}\mathcal{D}_{k}$ for better 3D reconstruction. In other words, we need to collect those datasets that could change the GS model from $\mathcal{S}^{{}^{\prime}}$ to $\mathcal{S}$ with the maximum extent. According to the uncertainty sampling theory [17, 29, 30], this is equivalent to maximizing the total inference loss of model $\mathcal{S}^{{}^{\prime}}$ on $\bigcup_{k\in\mathcal{X}}\mathcal{D}_{k}$, denoted as $C(\mathcal{S}^{{}^{\prime}},\{\mathcal{D}_{k}\},\mathbf{x})$. Specifically, the GS inference function $\mathcal{R}(\cdot|\mathcal{S}^{{}^{\prime}})$ is able to render a photo-realistic image from a camera pose as $\widehat{\mathbf{v}}_{i,k}=\mathcal{R}(\mathbf{s}_{i,k}|\mathcal{S}^{{}^{\prime}})$. To derive an explicit form of $C$, we need to measure the difference between the rendered image $\widehat{\mathbf{v}}_{i,k}$ (produced using $\mathcal{S}^{{}^{\prime}}$) and ground-truth image $\mathbf{v}_{i,k}$. According to [7], a proper metric is the vanilla GS that adopts the following loss function:
@@
respectively, with $\mathcal{L}_{\mathsf{SSIM}}$ being the SSIM function detailed in [42, Eqn. 5], and the weight $\lambda\in[0,1]$ is set to $\lambda=0.2$ according to [7].
@@ III Proposed Sample-Then-Transmit Strategy
The above $8$ steps can be categorized into two main stages, i.e., EGS loss prediction and EGS full transmission, where each stage consists of $4$ steps.
@@ IV EGS Loss Prediction
Given $\rho_{k}$, the next question is how to determine $\tilde{\mathcal{D}}_{k}$. A naive approach is to adopt random sampling, which randomly chooses an image inside $\mathcal{D}_{k}$ each time until $\lvert\tilde{\mathcal{D}}_{k}\rvert$ images are collected [43]. Another approach is to adopt uniform sampling, which orders the images in $\mathcal{D}_{k}$ according to the timestamps and selects an image for transmission for every $\lvert\mathcal{D}_{k}\rvert/\lvert\tilde{\mathcal{D}}_{k}\rvert$ image frames. These methods do not take the importance of images into account, and may lead to redundancy in transmitting similar images. For example, if the robot remains stationary, multiple images may be nearly identical, uploading them yields little additional value. To address the problem, the following subsection will present a method that improves sampling efficiency compared to the above approaches.